# DRAFT- Rotation Curve Fitting Model

### v2

@@ 6 Acknowledgments
This work is dedicated to Emmett Till. We acknowledge and express gratitude to the first nations peoples on whose unceded lands this was written; including but not limited to the the Coast Salish bands of the Puget Sound, the Cheyenne, Arapaho and Southern Ute Peoples of Colorado, and the Algonquian and Iroquoian Peoples of Massachusetts and New York, and the Navajo Peoples of New Mexico. The authors would like to thank V. P. Nair, T. Boyer, I. Chavel, R. Walterbos, N. P. Vogt, S. McGaugh, A. Klypin, R. Rivera, N. Oblath, J. Formaggio, J. Conrad, and P. Fisher.
| MW:SXJ | DM | DM | DM | RCFM | RCFM | RCFM | RAR | RAR | RAR | |

### v4

@@ 6 Acknowledgments
This work is dedicated to Emmett Till. We acknowledge and express gratitude to the first nations peoples on whose ceded and unceded lands this was written; including but not limited to the the Coast Salish bands of the Puget Sound, the Cheyenne and Arapaho Tribes and Ute Tribes of Colorado, the Algonquian and Iroquoian Peoples of Massachusetts and New York, and the Navajo and Pueblo Nations of New Mexico. The authors would like to thank V. P. Nair, T. Boyer, M. Kaku, I. Chavel, R. Walterbos, N. P. Vogt, S. McGaugh, A. Klypin, T. Quinn, S. Tuttle, M. Juric, R. Rivera, N. Oblath, J. Formaggio, J. Conrad, P. Fisher, Y. Sofue, C. Mihos, and M. Merrifield.
| MW:SXJ | Dark Matter | RCFM | RAR | |||||||
@@
Data Availability
The code used to generate all figures and fit results is publicly available at https://github.com/Cisneros-Galaxy/RCFM. The rotation curve data and photometric models used in this paper are from the SPARC database available at http://astroweb.cwru.edu/SPARC/. The few other datasets used as examples are available at the cited reference.